OpenStudy (anonymous):

fan & medal! A dilation has been performed on ΔABC. The length of AB is 4 inches and the length of A'B' is 24 inches. If segment BC has a length of 3 inches, what is the length of segment B'C'? 3 in. 18 in. 8 in. 0.5 in.

Vocaloid (vocaloid):

A'B' = 24 and AB = 4, that means that the scale factor used for the dilation is 6 because 24/4 is 6 B'C' = 6*BC

OpenStudy (anonymous):

so whats the answer ??????????????? I don't understand

Vocaloid (vocaloid):

BC = 3 B'C' = 6*BC